OpenStudy (anonymous):

what is the slope of the line passing through the points A(-5,7) andB(4,7)

OpenStudy (anonymous):

slope = difference in the y-values/difference in x-values

OpenStudy (anonymous):

ok so how do u do it

OpenStudy (anonymous):

what is the difference in y-values?

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

No. What are the two y-values?

OpenStudy (anonymous):

-5 and 4

OpenStudy (anonymous):

Those are the x-values. The y-values are 7 and 7. What is the difference?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

Yes. So you have a fraction with 0 in the numerator, so the entire fraction equals 0. So the slope is 0. And a line with a slope of 0 is always a horizontal line. If you plot the 2 points, you will see that it is a horizontal line (whose slope is 0).

OpenStudy (anonymous):

so all horazonal lines are 0 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Welcome. Now try doing this one: Find the slope of the line passing through the points (-2,4) and (3,-8).

OpenStudy (anonymous):

y= -4 x=-5

OpenStudy (anonymous):

You take the Difference between the two y-values: The two y-values are 4 and -8; the difference is -8-4 = -12 Take the difference between the two x-values: The 2 x-values are -2 and 3. The difference is 3 - (-2) = 3 + 2 = 5 Slope = -12/5.